What is the magnitude of the charge difference?
Answer #1

- 1142 Coulombs 4 e,r2 Law. Like charges repel Unlike charges attract 0 42

ー9x 109 N·m 2/C-Coulombs constant k = _

let the angle initially is \Theta = sin-1(0.0035/0.8) = 0.25

and charges q and Q

forces on each point charges are :

mg downwards

electrical force Fe, left for the left charge and right for the right charge

tension T in the string

so at equilibrium

mg = T cos\Theta

Fe = T sin\Theta

so Fe/mg = tan \Theta

kqQ/mgr2  = tan\Theta

kqQ/mg(0.0072) = 0.004375

qQ = 4.19*10-19 .......(1)

final charge on each is (q+Q)/2

and the angle \Theta' = 0.286

k(q+Q)2/4mgr2  = tan\Theta

k(q+Q)2/mg = 0.005*4*0.0082

(q+Q)2 = 25.01*10-19

q = 15.18*10-10 -Q ....(2)

from (1) and (2)

(15.18*10-10 -Q)Q = 4.19*10-19

Q2 -15.18*10-10Q +4.19*10-19 =0

Q = 12.21*10-10 or 3.7*10-10 C

so q = 15.18*10-10 -Q = 15.18*10-10 - 12.21*10-10 = 2.97*10-10 C

so the difference Q-q = (12.21-2.97)*10-10 = 9.24*10-10C = 0.92 nC